What is the molecular formula for the following compound? Select one: O a. C6H12 ob. C7H11 oc. C6H11 o d. C7H12 o e. C7H14
Answer #1

Number of Carbons = 6 from the ring + 1 from the methyl group = 6 + 1 = 7

Number of H = 3 from the methyl group + 1 + 2 * 3 + 2 * 1 = 12

Formula : C7H12